# Epicardium Prompt-guided Real-time Cardiac Ultrasound Frame-to-volume Registration

### v2

@@ Abstract
Real-time fusion of intraoperative 2D ultrasound images and preoperative 3D ultrasound volume based on the frame-to-volume registration can provide a comprehensive guidance view for cardiac interventional surgery. However, cardiac ultrasound images are characterized by a low signal-to-noise ratio and small differences between adjacent frames, coupled with significant dimension variations between 2D frames and 3D volumes to be registered, resulting in real-time and accurate cardiac ultrasound frame-to-volume registration being a very challenging task. This paper introduces a lightweight end-to-end Cardiac Ultrasound frame-to-volume Registration network, termed CU-Reg. Specifically, the proposed model leverages epicardium prompt-guided anatomical clues to reinforce the interaction of 2D sparse and 3D dense features, followed by a voxel-wise local-global aggregation of enhanced features, thereby boosting the cross-dimensional matching effectiveness of low-quality ultrasound modalities. We further embed an inter-frame discriminative regularization term within the hybrid supervised learning to increase the distinction between adjacent slices in the same ultrasound volume to ensure registration stability. Experimental results on the reprocessed CAMUS dataset demonstrate that our CU-Reg surpasses existing methods in terms of registration accuracy and efficiency, meeting the guidance requirements of clinical cardiac interventional surgery. Our code is available at https://github.com/LLEIHIT/CU-Reg.
@@ 4.0.1 Acknowledgements
This work was supported in part by the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No.: T45-401/22-N), in part by the Hong Kong Innovation and Technology Fund under Grant GHP/080/20SZ, in part by the Innovation and Technology Fund under Guangdong-Hong Kong Technology Cooperation Funding Scheme under Grant GHP/050/20SZ, in part by the National Natural Science Foundation of China under Grant 62273328, and in part by the Guangdong-Hong Kong-Macao Research Team Project under Grant 2021B1515130003.

### v3

@@ Abstract
Real-time fusion of intraoperative 2D ultrasound images and the preoperative 3D ultrasound volume based on the frame-to-volume registration can provide a comprehensive guidance view for cardiac interventional surgery. However, cardiac ultrasound images are characterized by a low signal-to-noise ratio and small differences between adjacent frames, coupled with significant dimension variations between 2D frames and 3D volumes to be registered, resulting in real-time and accurate cardiac ultrasound frame-to-volume registration being a very challenging task. This paper introduces a lightweight end-to-end Cardiac Ultrasound frame-to-volume Registration network, termed CU-Reg. Specifically, the proposed model leverages epicardium prompt-guided anatomical clues to reinforce the interaction of 2D sparse and 3D dense features, followed by a voxel-wise local-global aggregation of enhanced features, thereby boosting the cross-dimensional matching effectiveness of low-quality ultrasound modalities. We further embed an inter-frame discriminative regularization term within the hybrid supervised learning to increase the distinction between adjacent slices in the same ultrasound volume to ensure registration stability. Experimental results on the reprocessed CAMUS dataset demonstrate that our CU-Reg surpasses existing methods in terms of registration accuracy and efficiency, meeting the guidance requirements of clinical cardiac interventional surgery. Our code is available at https://github.com/LLEIHIT/CU-Reg.
@@ 4.0.1 Acknowledgements
This work was supported in part by the Research Grants Council of the Hong Kong Special Administrative Region under Grant T45-401/22-N, in part by the Hong Kong Innovation and Technology Fund under Grants GHP/080/20SZ and GHP/050/20SZ, in part by the National Natural Science Foundation of China under Grants U23A20391 and 62273328, and in part by the Regional Joint Fund of Guangdong under Grants 2021B1515130003 and 2021B1515120011.